# Worldline EFT treatment of quadratic and cubic gravity theories

### v1

@@ Abstract
This paper explores modifications to General Relativity (GR) by considering higher-order curvature terms in the gravitational action, specifically focusing on the quadratic Ricci scalar and a particular cubic contraction of the Riemann tensor. These modifications introduce new interactions at short distances, potentially altering the dynamics of compact objects. We calculate the effective two-body binding potential energy for these modified theories to quantify these effects using the worldline effective field theory (WEFT) formalism. This approach allows us to systematically integrate out short-distance gravitational effects, capturing the modifications to the binding potential. Our results demonstrate how the quadratic Ricci scalar and cubic Riemann tensor terms contribute to the two-body interaction at the leading order, highlighting deviations from classical GR predictions. These findings provide insight into the possible observational signatures of such modified gravity theories in binary systems and other astrophysical settings.
@@ I Introduction and Motivation
Ever since [1] developed a novel method for describing the dynamics of General Relativity (GR) using the Worldline Effective Field Theory (WEFT), interest in it has spiked. Using the path integral approach of Quantum Field Theory (QFT), one can now study both conservative and radiative dynamics of the binary problem in gravity. While the calculations of the binary problem have progressed quite a lot in the standard post-Newtonian (PN) formalism, it is slow and cumbersome. WEFT of GR has delivered results for the conservative and radiative dynamics in a few years, compared to what it took for the PN formalism to achieve in some decades. This efficiency has attracted a significant amount of attention to the field of WEFTs in recent years. Many reviews of the classical aspects of WEFTs have come up, which include [2, 3, 4, 5], as well as quantum gravity effects like [6, 7, 8, 9, 10, 11].
Since the WEFT approach to GR has been tackled to a very high PN order, the next logical step is modifying the gravitational action. Modified theories of gravity have gained a lot of traction since the very early days of GR, whenever it was evident that GR was not a complete theory of gravity. The singularity at the centre of a black hole (BH), the flattening of the rotation curves of galaxies, the cosmological constant problem, etc., have plagued GR. Two approaches to dealing with the problems of GR have emerged: string theory and loop quantum gravity. At the classical limit, both theories are expected to reproduce GR. This study will deal with the first relevant subdominant classical effect of a possible high-energy theory of gravity.
@@
where $R$ is the Ricci Scalar. In the renormalization program of GR, it was found by [12] that GR was renormalizable at one loop order. It was found that the divergences could be encoded in curvature squared terms like $R^{2},R_{\mu\nu}R^{\mu\nu},R_{\mu\alpha\rho\beta}R^{\mu\alpha\rho\beta}$. However, it was also found in [12] and by later authors that the curvature squared term could be field redefined away by a suitable metric redefinition. Hence, it is not prudent to include curvature squared terms for a proper theory of modified gravity in the absence of matter or BH spacetime. However, as a proof of concept, we will show the leading order change in the two-particle effective potential for the Ricci squared term added to the EH action.
It has been known for quite some time that the S matrix of GR is divergent (non-renormalizable) at two-loop order in four dimensions [13, 14]. The two loop divergences correspond to curvature cubed terms (scalars) and quadratic scalars of the covariant derivative of Ricci scalars and tensors. Utilizing the field equations makes most of the terms vanish except two Riemann cubed terms, of which only one is independent. Therefore, all two-loop divergences of the EH action can be parameterized by the following term
@@
While the particular curvature cubed term of Eq. (2) appears in the renormalization program as an actual divergence that cannot be renormalized, at a more recent time, a different procedure of field redefinition led to the same result. In [15], it was shown that field redefinitions of the metric tensor could get rid of all quadratic terms as well all but two Riemann curvature cubed terms from an action containing higher dimensional terms added to the EH action as an EFT. It was further shown in the similar spirit of [13, 14] by [16] that out of the two Riemann cubed terms, only one is independent and is again given by
@@
where $\lambda_{ev}$ is some constant parametrizing the strength of the term in the action. This Riemann cubed term of Eq. (3) cannot be field redefined away. Thus, two seemingly different approaches seem to lead to the same Riemann cubed term, implying a connection between the renormalization program and field redefinition strategies. Hence, when added to the EH action, the Riemann cubed term can be a candidate for the first subdominant effect of a low energy limit of a high energy theory of gravity in vacuum. Therefore, in this study, we find the two-particle effective potential, using the WEFT formalism, of the EH action plus the Riemann cubed term, that is
@@
where $\alpha$ here is a fundamental constant that quantifies the cubic term’s strength. Non-rotating black hole solutions for this theory were found in [15]. The slowly rotating solution (along with the cubic odd parity part) was found in [17], and perturbations of the rotating solution were studied in [18]. The Quasi-normal mode frequencies for the non-rotating case for this particular theory and the odd parity part have been recently found in [19]. Recently, Ref. [20] studied the degrees of freedom in this model and found that the theory only carries a spin-2 field but features a preferred frame at short distances, suggesting a possible connection to unknown ultraviolet physics.
@@ II Worldline EFT of GR: a mini review
Worldline effective field theories (EFTs) of gravity are a powerful framework used to study the dynamics of massive objects, such as black holes or neutron stars, in GR and modified theories of gravity. They combine quantum field theory and general relativity techniques to simplify the complex gravitational interactions between massive bodies. This particular framework has proven itself to be very efficient in obtaining various PN results of GR compared to the standard PN formalism, which required decades of research. The word ‘effective’ in EFTs of gravity or other QFTs implies that an approximation is being made when it comes to the scales (which can be energy, for example) of the problem. In theories which have phenomena covering a wide range of energy or length scales and where the various scales are decoupled from each other, EFTs are used to calculate the effects of a different energy (length) scale, usually higher, on the energy (length) scale in question, usually lower. In the case of gravity, short (long) distance effects are synonymous with high (low) energies.
The compact binary problem of gravity is one such problem where there is a clear separation of scales at the early in-spiral, or when the factor $\frac{v}{c}$ ($v$ being the characteristic orbital velocity) is close to zero (non-relativistic). In such a situation, the hierarchy of scales become
@@
In the hierarchy of scales, we first encounter the physics in the region of one of the compact objects we are interested in. A compact object, at the first approximation, can be thought of as a point particle whose centre of mass is moving along a worldline $x^{\mu}(\lambda)$ that is coupled to the gravitational field $g_{\mu\nu}$, where $\lambda$ is an arbitrary affine parameter acting as some time, or can be taken as the proper time along its trajectory. However, real compact objects in nature are not point particles but have some ‘spread’. This spread can be perturbatively treated as multipole moments of the source energy-momentum tensor. This requires an in-depth knowledge of the source energy-momentum tensor. However, we generally do not have complete information on the energy-momentum tensor. In such a case, we can parametrize our ignorance about the spread of the compact object as a series of operators with Wilson coefficients proportional to the so-called ‘Love numbers’ of the compact object. This spread and the Love numbers are directly related to the tidal response of one of the compact objects in the presence of the other. Thus, the effective point particle action can be written as follows
[…]
where in the first term $d\tau=\sqrt{g_{\mu\nu}dx^{\mu}dx^{\nu}}$, and is the result of the geodesic motion with respect to the metric $g_{\mu\nu}$. The second and the third term with Wilson coefficients $c_{S/T}$ can be removed by a field redefinition of the metric, as was argued in [1]. Therefore, they can be removed. The third and fourth terms contain contractions of the so-called electric and magnetic parts of the Riemann tensor defined as follows:
@@
However, since we are only interested in BBH space-times in this study, the constants $c_{E}$ and $c_{B}$ also vanish, as was shown in [21]. Hence, we are simply left with the first term, that is
[…]
@@
A step up from the smallest length scale $d$ is the near zone $r$. In this zone, the weak field regime applies, and one can expand the metric tensor about the flat spacetime $\eta_{\mu\nu}$ as
@@
To fully exploit the separation of scales in the binary problem, one usually splits further the perturbation $h_{\mu\nu}$ into potential $(H_{\mu\nu})$, corresponding to the short length scale, and radiation $(\bar{h}_{\mu\nu})$ modes, corresponding to the larger length scale, as was done in [3] and [5]. This separates the physics of $r$ from the physics of $\lambda$. The potential modes are treated as near instantaneous, given that the rate of change of the configuration of the binary is much slower than the rate at which gravitons are exchanged between the two BHs. To get the conservative dynamics, or the binding potential of the binary, one needs to integrate out both the potential and radiation mode, that is
@@
This paper, however, does not use the decomposition of the potential+radiation method of the metric perturbation; instead, it uses a Kaluza-Klein decomposition of the metric/metric perturbation. Given a d+1 dimensional spacetime, we can rewrite the metric components, as was done in [22], as follows
@@
We obtain the interaction terms for the perturbative path integral treatment by expanding the three terms of the action $(S_{EH},S_{pp},S_{GF})$ till the fifth order in $h_{\mu\nu}$ and then decompose $h$ into $\left(\phi,A_{i},\sigma_{jk}\right)$, as shown in Eq. (15). We also scale the fields $\left(\phi,A_{i},\sigma_{jk}\right)$ in powers of $c$ to obtain a PN series of the interaction terms in the bulk and worldline action in the following manner
@@
[…]
We have found that the 2 PN two-particle effective Lagrangian does not change after including the fifth order of perturbation of the metric in the action, as well as reproduces the same from previous literature [23], where they quote effective Lagrangian for only one of the particles. We reproduce the two-particle effective binding potential energy $V_{eff}$ (after summing over both the particles) till 2 PN in this paper for the benefit of the reader as follows:
@@
$V_{2PN}$ $=$ $-Gm_{1}m_{2}\left[-\frac{1}{8r^{3}}\left\{v_{1}^{i}\left[v_{1}^{j}x_{j}\left\{x_{k}\left(6v_{2}^{k}v_{1,i}-12v_{2}^{k}v_{2,i}-a_{2}^{k}x_{i}\right)+x_{i}v_{2}^{k}v_{2,k}\right\}+v_{2}^{j}v_{2}^{k}x_{k}\left(6x_{i}v_{2,j}+x_{j}v_{1,i}\right)\right]\right.\right.$ (23)
$\left.+a_{1}^{i}x_{i}v_{2}^{j}x_{j}v_{2}^{k}x_{k}\right\}+\frac{1}{8r}\left\{a_{1}^{i}x_{i}v_{2}^{j}v_{2,j}-14a_{1}^{i}v_{2}^{j}x_{j}v_{2,i}+v_{1}^{i}v_{1,i}\left(7v_{1}^{j}v_{1,j}-10v_{1}^{j}v_{2,j}+3v_{2}^{j}v_{2,j}\right.\right.$
@@
$\left.\left.-a_{1}^{i}a_{2}^{j}x_{i}x_{j}\right\}+\frac{15}{8}ra_{1}^{i}a_{2,i}+\frac{3v_{1}^{i}x_{i}v_{1}^{j}x_{j}v_{2}^{k}x_{k}v_{2}^{l}x_{l}}{8r^{5}}\right]-G^{2}m_{1}m_{2}\left[\frac{1}{4r^{2}}\left\{m_{1}\left[8v_{1}^{i}v_{1,i}+7\left(v_{2}^{i}-2v_{1}^{i}\right)v_{2,i}\right]\right.\right.$
@@
where $r=\sqrt{x^{i}x_{i}}$ is the distance between the two point particles, $x^{i}$ is the separation vector, $a^{i}_{K}$ for $K=1,2$ is the acceleration vector and $v^{i}_{K}$ for $K=1,2$ is the velocity vector for the first and the second object. We only give the effective potential and not the PN kinetic energy since it does not change under the change in the bulk action. In the rest of this study, we keep the worldline unchanged while we change the bulk action.

### v2

@@ Abstract
This paper explores modifications to General Relativity (GR) by considering higher-order curvature terms in the gravitational action, specifically focusing on the quadratic Ricci scalar and a particular cubic contraction of the Riemann tensor. These modifications introduce new interactions at short distances, potentially altering the dynamics of compact objects. We calculate the effective two-body binding potential energy for these modified theories to quantify these effects using the worldline effective field theory (WEFT) formalism. This approach allows us to systematically integrate out short-distance gravitational effects, capturing the modifications to the binding potential. Our results demonstrate how the quadratic Ricci scalar and cubic Riemann tensor terms contribute to the two-body interaction at the leading order, highlighting deviations from classical GR predictions. These findings offer insight into the potential observational signatures of modified gravity theories in binary systems and other astrophysical settings.
@@ I Introduction and Motivation
Ever since [1] developed a novel method for describing the dynamics of General Relativity (GR) using the Worldline Effective Field Theory (WEFT), interest in it has spiked. Using the path integral approach of Quantum Field Theory (QFT), one can now study both conservative and radiative dynamics of the binary problem in gravity. While the calculations of the binary problem have progressed significantly in the standard post-Newtonian (PN) formalism, they remain slow and cumbersome. WEFT of GR has delivered results for conservative and radiative dynamics in a few years, compared to what it took for the PN formalism to achieve over several decades. This efficiency has attracted a significant amount of attention to the field of WEFTs in recent years. Many reviews of the classical aspects of WEFTs have come up, which include [2, 3, 4, 5], as well as quantum gravity effects like [6, 7, 8, 9, 10, 11].
Since the WEFT approach to GR has been tackled to a very high PN order, the next logical step is modifying the gravitational action. Modified theories of gravity have gained significant traction since the early days of GR, when it became evident that GR was not a complete theory of gravity. The singularity at the centre of a black hole (BH), the flattening of rotation curves of galaxies, the cosmological constant problem, and others have plagued GR. Two approaches to dealing with the problems of GR have emerged: string theory and loop quantum gravity. At the classical limit, both theories are expected to reproduce GR. This study will deal with the first relevant subdominant classical effect of a possible high-energy theory of gravity.
@@
where $R$ is the Ricci Scalar. In the renormalization program of GR, it was found by [12] that GR was renormalizable at one loop order. It was found that the divergences could be encoded in curvature squared terms like $R^{2},R_{\mu\nu}R^{\mu\nu},R_{\mu\alpha\rho\beta}R^{\mu\alpha\rho\beta}$. However, it was also found in [12] and by later authors that the curvature squared term could be field redefined away by a suitable metric redefinition. Hence, it is not prudent to include curvature squared terms for a proper theory of modified gravity in the absence of matter or in BH spacetimes. However, as a proof of concept, we will show the leading order change in the two-particle effective potential for the Ricci squared term added to the EH action.
It has been known for quite some time that the S matrix of GR is divergent (non-renormalizable) at two-loop order in four dimensions [13, 14]. The two loop divergences correspond to curvature cubed terms (scalars) and quadratic scalars of the covariant derivative of Ricci scalars and tensors. Utilising the field equations makes most of the terms vanish except for two Riemann cubed terms, of which only one is independent. Therefore, all two-loop divergences of the EH action can be parametrized by the following term
@@
While the particular curvature cubed term of Eq. (2) appears in the renormalization program as an actual divergence that cannot be renormalized, at a more recent time, a different procedure of field redefinition led to the same result. In [15], it was shown that field redefinitions of the metric tensor could get rid of all quadratic terms, as well as all but two Riemann curvature cubed terms, from an action containing higher-derivative terms added to the EH action as an EFT. It was further shown in the similar spirit of [13, 14] by [16] that out of the two Riemann cubed terms, only one is independent and is again given by
@@
where $\lambda_{ev}$ is some constant parametrizing the strength of the term in the action. This Riemann cubed term of Eq. (3) cannot be field redefined away. Thus, two seemingly different approaches seem to lead to the same Riemann cubed term, implying a connection between the renormalization program and field redefinition strategies. Hence, when added to the EH action, the Riemann cubed term can be a candidate for the first subdominant effect of a low-energy limit of a high-energy theory of gravity in vacuum. Therefore, in this study, we find the two-particle effective potential, using the WEFT formalism, of the EH action plus the Riemann cubed term, that is
@@
where $\alpha$ here is a fundamental constant that quantifies the cubic term’s strength. Non-rotating black hole solutions for this theory were found in [15]. The slowly rotating solution (along with the cubic odd parity part) was found in [17], and perturbations of the rotating solution, as well as corrections to the quasinormal modes (to the first order in spin) were studied in [18]. The Quasi-normal mode frequencies for the non-spinning case, as well as the odd parity part, have recently been found in [19] for this particular theory. Following this, Ref. [20] found the corrections to the quasinormal mode overtones for the spinning geometry. Recently, Ref. [21] studied the degrees of freedom in this model and found that the theory only carries a spin-2 field but features a preferred frame at short distances, suggesting a possible connection to unknown ultraviolet physics.
A post-Minkowskian study using modern scattering amplitude techniques was performed on a variant of the action (4) in Refs. [22, 23] where it was argued that in non-vacuum space-times, there exists another independent combination of Riemann cubed terms which can contribute to the effective potential. However, as was mentioned in [22], the three and four point graviton amplitudes of the extra combination vanishes. However, it was also argued that the extra combination can affect tidal terms at 5 PN order. However, in this study, we consider only BH space-times and therefore will not consider the additional combination. The above Refs. calculated the conservative dynamics of a compact binary system. Radiative dynamics of the variant of the action (4) was calculated in Ref. [24]. Finally, in Ref. [25], constraints were placed on the coefficient $\alpha$ and the coefficient of the extra combination using the GW170608 event.
@@ II Worldline EFT of GR: a mini review
Worldline effective field theories (EFTs) of gravity are a powerful framework used to study the dynamics of massive objects, such as black holes or neutron stars, in GR and modified theories of gravity. They combine quantum field theory and general relativity techniques to simplify the complex gravitational interactions between massive bodies. This particular framework has proven to be highly efficient in obtaining various PN results for GR compared to the standard PN formalism, which required decades of research. The word ‘effective’ in EFTs of gravity or other QFTs implies that an approximation is being made when it comes to the scales (which can be energy, for example) of the problem. In theories which have phenomena covering a wide range of energy or length scales and where the various scales are decoupled from each other, EFTs are used to calculate the effects of a different energy (length) scale, usually higher, on the energy (length) scale in question, usually lower. In the case of gravity, short (long) distance effects are synonymous with high (low) energies.
The compact binary problem of gravity is one such problem where there is a clear separation of scales at the early in-spiral, or when the factor $\frac{v}{c}$ ($v$ being the characteristic orbital velocity) is close to zero (non-relativistic). In such a situation, the hierarchy of scales becomes
@@
In the hierarchy of scales, we first encounter physics in the region of one of the compact objects we are interested in. A compact object, at the first approximation, can be thought of as a point particle whose centre of mass is moving along a worldline $x^{\mu}(\lambda)$ that is coupled to the gravitational field $g_{\mu\nu}$, where $\lambda$ is an arbitrary affine parameter acting as some time, or can be taken as the proper time along its trajectory. However, real compact objects in nature are not point particles but have some ‘spread’. This spread can be perturbatively treated as multipole moments of the source energy-momentum tensor. This requires an in-depth knowledge of the source energy-momentum tensor. However, we generally do not have complete information on the energy-momentum tensor. In such a case, we can parametrize our ignorance about the spread of the compact object as a series of operators with Wilson coefficients proportional to the so-called ‘Love numbers’ of the compact object. This spread and the Love numbers are directly related to the tidal response of one of the compact objects in the presence of the other. Thus, the effective point particle action can be written as follows
$S_{pp}=-mc^{2}\int d\tau+c_{S}\int d\tau R+c_{T}\int d\tau R_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}+c_{E}\int d\tau E_{\mu\nu}E^{\mu\nu}+c_{B}\int d\tau B_{\mu\nu}B^{\mu\nu}+\cdots$ (6)
where in the first term $d\tau=\sqrt{g_{\mu\nu}dx^{\mu}dx^{\nu}}$, and is the result of the geodesic motion with respect to the metric $g_{\mu\nu}$. The second and third terms with Wilson coefficients $c_{S/T}$ can be removed by a field redefinition of the metric, as argued in [1]. Therefore, they can be ignored. The fourth and fifth terms contain contractions of the so-called electric and magnetic parts of the Riemann tensor, defined as follows:
@@
However, since we are only interested in BBH space-times in this study, the constants $c_{E}$ and $c_{B}$ also vanish in GR, as was first shown in Refs. [26, 27] and was reconfirmed in Ref. [28] using scattering amplitude techniques. However, this is not in general true for modified gravity theories, especially for the cubic Riemann theory that we consider in this study, as were shown in Refs. [29, 30]. But it should also be noted that tidal terms are PN suppressed, and that at the leading order contributions due to the cubic Riemann term will be only due to the point particle minimal coupling to gravity, which is what we consider in this study. Hence, we are simply left with the first term, that is
$S_{pp}=-mc^{2}\int d\Lambda d^{4}x\delta^{4}\left[x^{\mu}-x^{\mu}_{pp}(\Lambda)\right]\sqrt{g_{\mu\nu}v^{\mu}v^{\nu}}$ (8)
@@
A step up from the smallest length scale, $d$, is the near zone, $r$. In this zone, the weak field regime applies, and one can expand the metric tensor about the flat spacetime $\eta_{\mu\nu}$ as
@@
To fully exploit the separation of scales in the binary problem, one usually splits further the perturbation $h_{\mu\nu}$ into potential $(H_{\mu\nu})$, corresponding to the short length scale, and radiation $(\bar{h}_{\mu\nu})$ modes, corresponding to the larger length scale, as was done in Ref. [1] and was further explored in the reviews [3, 5]. This separates the physics of $r$ from the physics of $\lambda$. The potential modes are treated as near instantaneous, given that the rate of change of the configuration of the binary is much slower than the rate at which gravitons are exchanged between the two BHs. To get the conservative dynamics, or the binding potential of the binary, one needs to integrate out both the potential and radiation mode, that is
@@
This paper, however, uses a Kaluza-Klein decomposition of the potential modes putting the radiation modes to zero since we are dealing only with the conservative part of the problem. Given a d+1 dimensional spacetime, we can rewrite the metric components, as was done in [31], as follows
@@
We obtain the interaction terms for the perturbative path integral treatment by expanding the three terms of the action $(S_{EH},S_{pp},S_{GF})$ till the sixth order in $h_{\mu\nu}$ and then decompose $h$ into $\left(\phi,A_{i},\sigma_{jk}\right)$, as shown in Eq. (15). We also scale the fields $\left(\phi,A_{i},\sigma_{jk}\right)$ in powers of $c$ to obtain a PN series of the interaction terms in the bulk and worldline action in the following manner
@@
The rationale behind this particular type of scaling is a lesson that is learnt from the PN expansion of GR. It is well known (for example, see Ref. [32]) that the time-time component of the metric perturbation in PN theory scale as $\frac{1}{c^{2}}$, the time-space component as $\frac{1}{c^{3}}$, and the space-space component as $\frac{1}{c^{4}}$. Hence, we scale the corresponding Kaluza-Klein fields in the same manner.
@@
We begin by drawing all possible Feynman diagrams, carefully accounting for the scaling and order of the vertices. The total amplitude is the sum of all the topologically distinct and connected diagrams. After evaluating the amplitude and by performing the momentum integrals for the internal lines, the inverse Fourier transform of the amplitude (after putting $d=3$) leads to a time integral $\int\mathcal{L}_{eff}\,dt$ from which we extract the effective Lagrangian $\mathcal{L}_{eff}$, and eventually the effective potential $V_{eff}$. The full effective Lagrangian at various PN orders till 2 PN, including the kinetic terms, have been given in Appendix B. Since we do not change the worldline action due a change in the theory of gravity, the kinetic terms remain the same as in GR.
We find that after including the sixth order of metric perturbation, the Kaluza-Klein expanded action stops changing, implying that inclusion of the sixth order of metric perturbation is enough to obtain the correct metric expanded action till the second PN order at least. Similarly, we have found that the 2 PN two-particle effective Lagrangian also does not change after including the sixth order of perturbation of the metric in the action, and reproduces the same as in previous literature in Refs. [33, 34], where they quote effective Lagrangian. We reproduce the two-particle effective binding potential energy $V_{eff}$ till 2 PN in this paper for the benefit of the reader as follows:
@@
[…]
@@
[…]
@@
where $r=\sqrt{x^{i}x_{i}}$ is the distance between the two point particles, $x^{i}$ is the separation vector, $a^{i}_{K}$ for $K=1,2$ is the acceleration vector and $v^{i}_{K}$ for $K=1,2$ is the velocity vector for the first and the second object. We only provide the effective potential and not the PN kinetic energy, since it remains unchanged under changes in the bulk action.